# Evolving the Euler rotation axis as a dynamical system, using the Euler vector and generalizations

### v1

@@ Abstract.
${{\bf{M}}}(t)={\boldsymbol{\omega}}^{(i)}(t_{1})({h^{i+1}}/(i+1)!)M_{0}(t_{1})/2+o(h^{i+1})$, from the proof of Lemma 6. So ${{\bf{M}}}(t)/|{{\bf{M}}}(t)|\\ =({\boldsymbol{\omega}}^{(i)}(t_{1})/|{\boldsymbol{\omega}}^{(i)}(t_{1})|)M_{0}(t_{1})sgn(h^{i+1})+o(1)$. Now $\theta(t)$ has already been chosen for $t_{0}\leq t\leq t_{1}$, and is to be chosen for $t_{1}<t\leq t_{2}$. For for $t_{0}<t<t_{2}$ and $t\neq t_{1}$, we need ${{\bf{M}}}(t)={{\bf{n}}}(t)\sin(\theta(t)/2)$, so $|{{\bf{M}}}(t)|=|\sin(\theta(t)/2)|$, and ${{\bf{n}}}(t)={{\bf{M}}}(t)/\sin(\theta(t)/2)=({{\bf{M}}}(t)/|{{\bf{M}}}(t)|)sgn(\sin(\theta(t)/2))$$\ =({\boldsymbol{\omega}}^{(i)}(t_{1})/|{\boldsymbol{\omega}}^{(i)}(t_{1})|)M_{0}(t_{1})sgn\{h^{i+1}\sin(\theta(t)/2)\}+o(1)$. In order that ${\bf{n}}(t)$ have the same limit from the right and left as $t\rightarrow t_{1}$, $\sin(\theta(t)/2)$ must change sign as $t$ crosses $t_{1}$ if $i$ is even, and not if $i$ is odd. Note $\theta(t_{1})$ must be either $2\pi l$ or $2\pi(l+1)$. If $i$ is even, continuity of $\theta$ then requires that, for $t_{1}<t\leq t_{2}$, it be chosen with $2\pi(l-1)\leq\theta(t)\leq 2\pi l$ if $\theta(t_{1})=2\pi l$, and $2\pi(l+1)\leq\theta(t)\leq 2\pi(l+2)$ if $\theta(t_{1})=2\pi(l+1)$. If $i$ is odd, $\theta$ must not change sign, so continuity requires $2\pi(l+1)\leq\theta(t)\leq 2\pi(l+2)$ for $t_{1}<t\leq t_{2}$.
@@
The proof of existence of $\dot{\bf{n}}$ at $t_{1}$ is similar, but much longer and more tedious. We’ll leave out some detail. By the stronger assumption, ${\boldsymbol{\omega}}(t)={\boldsymbol{\omega}}^{(i)}(t_{1}){h^{i}}/i!+{\boldsymbol{\omega}}^{(i+1)}(t_{1}){h^{i+1}}/(i+1)!+o(h^{i+1})$. From earlier, ${{\bf{M}}}(t)={\boldsymbol{\omega}}^{(i)}(t_{1})({h^{i+1}}/(i+1)!)M_{0}(t_{1})/2+o(h^{i+1})$. Since ${\boldsymbol{\omega}}^{(i)}(t_{1})\times{\boldsymbol{\omega}}^{(i)}(t_{1})=0$, we get ${\boldsymbol{\omega}}(t)\times{{\bf{M}}}(t)=o(h^{2i+1}$. Using this and $M_{0}(t)=M_{0}(t_{1})+o(h)$ in (9),$\dot{{\bf{M}}}(t)=$ ${\boldsymbol{\omega}}^{(i)}(t_{1})({h^{i}}/i!)M_{0}(t_{1})/2+{\boldsymbol{\omega}}^{(i+1)}(t_{1})({h^{i+1}}/(i+1)!)M_{0}(t_{1})/2+o(h^{i+1})$. Integrating, ${\bf{M}}(t)={\boldsymbol{\omega}}^{(i)}(t_{1})({h^{i+1}}/(i+1)!)M_{0}(t_{1})/2\\ +{\boldsymbol{\omega}}^{(i+1)}(t_{1})({h^{i+2}}/(i+2)!)M_{0}(t_{1})/2+o(h^{i+2})$.
@@
Proof of Theorem 4.
This follows by induction from Lemmas 5 and 6. Suppose $t_{0}<t_{1}<...<t_{n}$ satisfy ${\bf{M}}(t_{j})={\bf{0}}$ for $j=1,...,n$, and ${\bf{M}}(t)\neq{\bf{0}}$ for $t_{0}<t<t_{n}$ and $t\neq t_{j}$ for $j=1,...,n$. Assume ${\bf{n}}(t)$ and $\theta(t)$ have been defined continuously for $t_{0}\leq t\leq t_{n}$. Let $t_{n+1}=\inf\{t>t_{n}:{\bf{M}}(t)={\bf{0}}\}$, where $t_{n+1}=\infty$ if no such $t$ exists. By Lemma 5, $t_{n+1}>t_{n}$. By Lemma 6, ${\bf{n}}(t)$ and $\theta(t)$ can be continuously extended to $t_{0}\leq t\leq t_{n+1}$. By Lemma 5, there can be no cluster point of the $t_{n}$, so "Zeno’s paradox" cannot occur; that is, $t_{n}\rightarrow\infty$, and ${\bf{n}}(t)$ and $\theta(t)$ are defined for all $t$. ∎
Corollary 7 (Continuation of ${\bf{E}}$ at boundaries).
Let ${\boldsymbol{\omega}}$ satisfy the hypotheses of Theorem 4. Then there exists ${\bf{E}}(t)$ which is continuous for all $t\geq t_{0}$, and satisfies Equation (7) except where $|{\bf{E}}(t)|=|\theta(t)|$ is a positive multiple of $2\pi$ and the r.h.s of the equation is not defined. If ${\boldsymbol{\omega}}$ satisfies the stronger differentiability condition of the theorem, ${\bf{E}}(t)$ is differentiable for all $t\geq t_{0}$
@@
Non-trivial example of going through $2\pi$.
@@
We’ll use the endpoint of this first run to create initial conditions for a second run. Let ${\boldsymbol{\omega^{*}}}(t)=-{\boldsymbol{\omega}}(t_{1}-t)={-\bf{i}}-(t_{1}-t){\bf{j}},t\geq 0$. Let $\theta^{*}(0)=2\pi-\theta_{1}$, ${\bf{n}}^{*}(0)=-{\bf{n}}_{1}$, and ${\bf{E}}^{*}(0)={\bf{n}}^{*}(0)\theta^{*}(0)$. Now run Equation (7) for ${t\geq 0}$, using ${\boldsymbol{\omega^{*}}}$ and ${\bf{E}}^{*}$ with these initial conditions. It can be shown that then $\theta^{*}(t)=2\pi-\theta(t_{1}-t)$ and ${\bf{n}}^{*}(t)=-{\bf{n}}(t_{1}-t)$ for $0\leq t\leq t_{1}$, where $\theta^{*}(t)=|{\bf{E}}^{*}(t)|$ and ${\bf{n}}^{*}(t)={\bf{E}}^{*}(t)/|{\bf{E}}^{*}(t)|$. Now $\theta(t_{1}-t)$ approaches zero as $t\rightarrow t_{1}^{-}$, so $\theta^{*}(t)$ approaches $2\pi$, and we have our example. Since ${\boldsymbol{\omega^{*}}}(t_{1})\neq{\bf{0}}$, $\theta^{*}(t)$ should go past $2\pi$ as $t$ passes $t_{1}$, according to Theorem 4.

### v2

@@ Abstract.
${{\bf{M}}}(t)={\boldsymbol{\omega}}^{(i)}(t_{1})({h^{i+1}}/(i+1)!)M_{0}(t_{1})/2+o(h^{i+1})$, from the proof of Lemma 5.3. So
${{\bf{M}}}(t)/|{{\bf{M}}}(t)|=({\boldsymbol{\omega}}^{(i)}(t_{1})/|{\boldsymbol{\omega}}^{(i)}(t_{1})|)M_{0}(t_{1})sgn(h^{i+1})+o(1)$. Now $\theta(t)$ has already been chosen for $t_{0}\leq t\leq t_{1}$, and is to be chosen for $t_{1}<t\leq t_{2}$. For for $t_{0}<t<t_{2}$ and $t\neq t_{1}$, we need ${{\bf{M}}}(t)={{\bf{n}}}(t)\sin(\theta(t)/2)$, so $|{{\bf{M}}}(t)|=|\sin(\theta(t)/2)|$, and ${{\bf{n}}}(t)={{\bf{M}}}(t)/\sin(\theta(t)/2)=({{\bf{M}}}(t)/|{{\bf{M}}}(t)|)sgn(\sin(\theta(t)/2))$$\ =({\boldsymbol{\omega}}^{(i)}(t_{1})/|{\boldsymbol{\omega}}^{(i)}(t_{1})|)M_{0}(t_{1})sgn\{h^{i+1}\sin(\theta(t)/2)\}+o(1)$. In order that ${\bf{n}}(t)$ have the same limit from the right and left as $t\rightarrow t_{1}$, $\sin(\theta(t)/2)$ must change sign as $t$ crosses $t_{1}$ if $i$ is even, and not if $i$ is odd. Note $\theta(t_{1})$ must be either $2\pi l$ or $2\pi(l+1)$. If $i$ is even, continuity of $\theta$ then requires that, for $t_{1}<t\leq t_{2}$, it be chosen with $2\pi(l-1)\leq\theta(t)\leq 2\pi l$ if $\theta(t_{1})=2\pi l$, and $2\pi(l+1)\leq\theta(t)\leq 2\pi(l+2)$ if $\theta(t_{1})=2\pi(l+1)$. If $i$ is odd, $\theta$ must not change sign, so continuity requires $2\pi(l+1)\leq\theta(t)\leq 2\pi(l+2)$ for $t_{1}<t\leq t_{2}$.
@@
The proof of existence of $\dot{\bf{n}}$ at $t_{1}$ is similar, but much longer and more tedious. We’ll leave out some detail. By the stronger assumption, ${\boldsymbol{\omega}}(t)={\boldsymbol{\omega}}^{(i)}(t_{1}){h^{i}}/i!+{\boldsymbol{\omega}}^{(i+1)}(t_{1}){h^{i+1}}/(i+1)!+o(h^{i+1})$. From earlier, ${{\bf{M}}}(t)={\boldsymbol{\omega}}^{(i)}(t_{1})({h^{i+1}}/(i+1)!)M_{0}(t_{1})/2+o(h^{i+1})$. Since ${\boldsymbol{\omega}}^{(i)}(t_{1})\times{\boldsymbol{\omega}}^{(i)}(t_{1})=0$, we get ${\boldsymbol{\omega}}(t)\times{{\bf{M}}}(t)=o(h^{2i+1}$. Using this and $M_{0}(t)=M_{0}(t_{1})+o(h)$ in (9),$\dot{{\bf{M}}}(t)=$ ${\boldsymbol{\omega}}^{(i)}(t_{1})({h^{i}}/i!)M_{0}(t_{1})/2+{\boldsymbol{\omega}}^{(i+1)}(t_{1})({h^{i+1}}/(i+1)!)M_{0}(t_{1})/2+o(h^{i+1})$. Integrating,
${\bf{M}}(t)={\boldsymbol{\omega}}^{(i)}(t_{1})({h^{i+1}}/(i+1)!)M_{0}(t_{1})/2+{\boldsymbol{\omega}}^{(i+1)}(t_{1})({h^{i+2}}/(i+2)!)M_{0}(t_{1})/2+o(h^{i+2})$.
@@
Proof of Theorem 5.1.
This follows by induction from Lemmas 5.2 and 5.3. Suppose $t_{0}<t_{1}<...<t_{n}$ satisfy ${\bf{M}}(t_{j})={\bf{0}}$ for $j=1,...,n$, and ${\bf{M}}(t)\neq{\bf{0}}$ for $t_{0}<t<t_{n}$ and $t\neq t_{j}$ for $j=1,...,n$. Assume ${\bf{n}}(t)$ and $\theta(t)$ have been defined continuously for $t_{0}\leq t\leq t_{n}$. Let $t_{n+1}=\inf\{t>t_{n}:{\bf{M}}(t)={\bf{0}}\}$, where $t_{n+1}=\infty$ if no such $t$ exists. By Lemma 5.2, $t_{n+1}>t_{n}$. By Lemma 5.3, ${\bf{n}}(t)$ and $\theta(t)$ can be continuously extended to $t_{0}\leq t\leq t_{n+1}$. By Lemma 5.2, there can be no cluster point of the $t_{n}$, so "Zeno’s paradox" cannot occur; that is, $t_{n}\rightarrow\infty$, and ${\bf{n}}(t)$ and $\theta(t)$ are defined for all $t$. ∎
Corollary 5.4 (Continuation of ${\bf{E}}$ at boundaries).
Let ${\boldsymbol{\omega}}$ satisfy the hypotheses of Theorem 5.1. Then there exists ${\bf{E}}(t)$ which is continuous for all $t\geq t_{0}$, and satisfies Equation (7) except where $|{\bf{E}}(t)|=|\theta(t)|$ is a positive multiple of $2\pi$ and the r.h.s of the equation is not defined. If ${\boldsymbol{\omega}}$ satisfies the stronger differentiability condition of the theorem, ${\bf{E}}(t)$ is differentiable for all $t\geq t_{0}$
@@
Example (Non-trivially going through $2\pi$).
@@
We’ll use the endpoint of this first run to create initial conditions for a second run. Let ${\boldsymbol{\omega^{*}}}(t)=-{\boldsymbol{\omega}}(t_{1}-t)={-\bf{i}}-(t_{1}-t){\bf{j}},t\geq 0$. Let $\theta^{*}(0)=2\pi-\theta_{1}$, ${\bf{n}}^{*}(0)=-{\bf{n}}_{1}$, and ${\bf{E}}^{*}(0)={\bf{n}}^{*}(0)\theta^{*}(0)$. Now run Equation (7) for ${t\geq 0}$, using ${\boldsymbol{\omega^{*}}}$ and ${\bf{E}}^{*}$ with these initial conditions. It can be shown that then $\theta^{*}(t)=2\pi-\theta(t_{1}-t)$ and ${\bf{n}}^{*}(t)=-{\bf{n}}(t_{1}-t)$ for $0\leq t\leq t_{1}$, where $\theta^{*}(t)=|{\bf{E}}^{*}(t)|$ and ${\bf{n}}^{*}(t)={\bf{E}}^{*}(t)/|{\bf{E}}^{*}(t)|$. Now $\theta(t_{1}-t)$ approaches zero as $t\rightarrow t_{1}^{-}$, so $\theta^{*}(t)$ approaches $2\pi$, and we have our example. Since ${\boldsymbol{\omega^{*}}}(t_{1})\neq{\bf{0}}$, $\theta^{*}(t)$ should go past $2\pi$ as $t$ passes $t_{1}$, according to Theorem 5.1.